OpenStudy (anonymous):

\[n - \sqrt{n ^{2} - n} \] Covergent or Divergent?

OpenStudy (anonymous):

Convergent. Converges to 0: \[n-\sqrt{n(n-1)}\] For extremely large values of n, the difference between n and n-1 is negligible, so basically you have: \[\sqrt{n(n-1)} \approx \sqrt{n(n)} = \sqrt{n^{2}} = n\] Thus you have n - n = 0